# Degeneration of genus three Ceresa cycles and limit of archimedean height

### 02 Apr 2026 v1

@@ Abstract.
For a family of smooth, projective curves of genus three, we interpret the limit of archimedean height pairings of Ceresa cycles associated with the family in terms of the archimedean pairing of algebraic cycles naturally associated with the central fiber of the family.
@@
S. Goswami was partially supported by the Universitat de Barcelona María Zambrano postdoctoral fellowship, and the Spanish MINECO research project PID2023-147642NB-I00. I. Spelta is a member of GNSAGA (INdAM) and partially supported by the Spanish MINECO research project PID2023-147642NB-I00.
@@ Introduction
In this article, we want to consider a family $\mathcal{C}\to\Delta$ of projective curves of fixed genus $g\geq 3$, with smooth general fiber $C_{t},\penalty 10000\ t\neq 0$, and central fiber $C_{0}$ having nodal singularities. Two functions can naturally be attached to such a family. The first one is a holomorphic function $\nu(t)\coloneqq AJ(Z(C_{t}))$, known as a normal function and defined using the Abel-Jacobi morphism. Both $\nu$ and its derivative, the Griffiths infinitesimal invariant, has been studied widely by several authors ([Gr83], [Voi88], [Hai90], [CP95],[KRP03], [GGK10], and [Hai25] to mention only a few). The second function is the height, which we introduce more generally. Let $X$ be a smooth complex projective variety, and let $Z$ and $W$ be homologically trivial algebraic cycles of codimensions $p$ and $q$, respectively, intersecting properly and satisfying $\dim X+1=p+q$. Under these assumptions, one necessarily has $|Z|\cap|W|=\emptyset$. To such a pair, one can associate a unique real number $\operatorname{ht}(Z,W)$, called the archimedean height. Hodge theoretically, $\operatorname{ht}(Z,W)$ measures how far a canonical subquotient $B_{Z,W}$ of the (rational) mixed Hodge structure $H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))$ is from being $\mathds{R}$-split. $B_{Z,W}$ is an example of a biextension mixed Hodge structure with graded pieces
@@
A natural problem is to extend this construction from $\mathcal{M}_{g}$ to its Deligne–Mumford compactification $\overline{\mathcal{M}}_{g}$. It can be shown that while the biextension line bundle extends algebraically, the metric does not (see [Lear90], [HR04], [Pe06, Section 5]). In the language of heights, for a family of biextensions $\{B_{t}\}_{t\neq 0}$, instead of $\operatorname{ht}(B_{t})$, the function $\operatorname{ht}(B_{t})+\mu\log|t|$ extends continuously to $t=0$, for $\mu\in\mathds{Q}$. The number $\mu$ is called the height jump, and it measures the singular behaviour of the height near the boundary. Subsequently, in [BP19], the authors studied this phenomenon in a very general setting, relating height jumps of variations of abstract biextension mixed Hodge structures over a complex manifold to canonical pairings in intersection cohomology.
While the articles above study the phenomenon of height jump in great detail, the aim of our paper is slightly different. Heuristically, we want to connect a certain limit height, which was first introduced in [BP19, Definition 70], to the archimedean height of a pair of cycles that ‘is defined on the boundary’. As the first non-trivial case, and following the approaches of [Co97, Hai90, KRP03], we consider the limit of heights of Ceresa cycles associated to genus three curves. We summarize our results in the following theorem:
@@
Let $\mathcal{C}\to\Delta$ be a family of smooth projective curves of fixed genus $g=3$ degenerating to the central fiber $C_{0}$, which is one of the following two types:
@@
    $C_{0}=C_{1}\cup C_{2}$, for two irreducible and smooth components $C_{1}$ and $C_{2}$ of genus one and two, respectively, with a separating node $y_{0}=C_{1}\cap C_{2}$.
Then, fixing a choice of coordinate $t\in\Delta$ such that the limit mixed Hodge structure $H^{1}_{\lim}(C_{t};\mathds{Z})$ is split (assumption 2.2), we can find a family of Ceresa cycles $\{Z(C_{t}),W(C_{t})\}_{t\in\Delta^{\ast}}$ that intersects properly in $JC_{t}$, such that the cycles $Z(C_{0})$ and $W(C_{0})$ also intersect properly in the compactification $X_{0}$ of the Jacobian $JC_{0}$, and are homologous to zero (Propositions 4.1 and 4.5). Finally, one has
$\lim_{t\to 0}\operatorname{ht}(Z(C_{t}),W(C_{t}))=\operatorname{ht}(\widetilde{Z}(C_{0}),\widetilde{W}(C_{0})),$
where $\widetilde{Z}(C_{0})$ and $\widetilde{W}(C_{0})$ are the lifts of $Z(C_{0})$ and $W(C_{0})$ to the normalization $\widetilde{X}_{0}$, and are naturally homologous to zero.
We now give an idea of the proof. The main part concerns the case of an irreducible nodal central fiber. In this situation, there is a nontrivial vanishing cycle and a monodromy operator $N\neq 0$ with $N^{2}=0$. The case where $C_{0}$ is reducible is much simpler, since $N=0$. Our argument is inspired by the work in [KRP03] and [BP19], and can be thought of as a bridge connecting the two. For a biextension variation $B_{Z_{t},W_{t}},\penalty 10000\ t\in\Delta^{\ast}$ arising for a family of pairs of null homologous cycles $\{Z_{t},W_{t}\}$ , [BP19, Definition 70] defines a limit height $H(N,F_{\infty},W)$ such that
$\lim_{t\to 0}(\operatorname{ht}(Z_{t},W_{t})+\mu\log|t|)=H(N,F_{\infty},W),$
where $\mu\in\mathds{Q}$, as mentioned before, is the height jump. In the cases of our interest, the height jump $\mu=0$. Hence, the key point is to show that $H(N,F_{\infty},W)=\operatorname{ht}(\widetilde{Z}(C_{0}),\widetilde{W}(C_{0}))$. As a first step, we have to find the limit of the biextension variation
@@
The assumptions 1.8 and 2.2 make this study significantly easier. Nevertheless, the resulting limit mixed Hodge structure $\mathcal{V}_{\lim}$ no longer has the structure of a biextension. Consequently, the analysis of the limiting height amounts to extracting from $\mathcal{V}_{\lim}$ the graded pieces of the biextension $H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}(C_{0})|,|\widetilde{W}(C_{0})|;\mathds{Q}(2))$. This is the core motivation of Theorem 4.7, and is achieved using a purity result for the cohomology of ‘mildly’ singular varieties (see Proposition 4.3), and the injectivity of the Clemens-Schmidt morphism in our situation. Finally, the equality of the two heights is achieved using an argument similar to Proposition 3.4.
We end with the following observation: Since by assumption pqp-q is torsion in $J\widetilde{C}_{0}$, we can define a covering map $\phi\colon\widetilde{X}_{0}\to J\widetilde{C}_{0}\times\mathds{P}^{1}$, and in turn a pair of higher cycles $\phi(\widetilde{Z}(C_{0}))=Z(\widetilde{C}_{0};p,q)$ and $\phi(\widetilde{W}(C_{0}))=W(\widetilde{C}_{0};p,q)$ in $K_{1}(J\widetilde{C}_{0})$ which are called Collino cycles. In [BGP22] and [BGP25], the first author defined a notion of height pairing of higher cycles on smooth projective varieties $X$ over $\mathds{C}$. The origin of this paper is linked to the natural question of comparing the limit height $H(N,F_{\infty},W)$ with the higher height of Collino cycles. However, it seems not to be the case, at least in the situation of our interest. This is unlike the normal function scenario, where regulators of higher cycles often arise naturally as limits of Abel-Jacobi normal functions (see, for example, [GGK10]).
The paper is organized as follows. In Section 1, we review basic material on mixed Hodge structures and on the Jacobian of a one-nodal curve, and we introduce the two types of degenerations used throughout the paper. In Section 2, we recall the theory of limit mixed Hodge structures of geometric origin and apply it to analyze the mixed Hodge structure $H^{3}_{\lim}(JC_{t};\mathds{Q})$. In Section 3, we discuss biextensions, both abstractly and in the geometric setting, introduce the notion of height, and recall the variational aspects of [BP19]. Finally, in Section 4, we study families of Ceresa cycles, establish their properties and prove Theorem 4.7.
@@ 1. Preliminaries
In this, section we gather preliminary results from Hodge theory, nodal curves, and Jacobians.
@@ 1.2. Deligne splitting
Next, we show that $\delta$ commutes with respect to a morphism of MHS.
@@ 1.3.1. Jacobian of a nodal curve
The semi-abelian variety $JC_{0}$ can be compactified (see [AK90, Mu07]). Indeed, identifying ${\rm Pic}^{0}(\widetilde{C}_{0})\cong{\rm Pic}^{0}(J\widetilde{C}_{0})$, the line bundle $L:=\mathcal{O}_{\widetilde{C}_{0}}(p-q)$ can be viewed as a line bundle over $J\widetilde{C}_{0}$, and $JC_{0}$ is the total space of $L$ minus the 0-section. Let $Y_{0}$ and $Y_{\infty}$ be the zero and infinity sections of $\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L)$ obtained through the projections $s_{0}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to\mathcal{O}_{J\widetilde{C}_{0}}$ and $s_{1}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to L$ respectively. Then, the compactification $X_{0}$ is obtained as
@@
in other words the singular locus of $X_{0}$ is given by $J\widetilde{C}_{0}$, which is smooth and projective. Elements in $X_{0}$ are rank one torsion-free sheaves $M$ of degree 0 whose sections $s$ are allowed to approach the point $x_{0}$. The normalization $\widetilde{X}_{0}$ is given by the projective bundle $\mathds{P}\left(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\right)$ over $J\widetilde{C}_{0}$. From [UBAP24, Proposition 5.3] and projective bundle formula for cohomology, the singular cohomology groups of $X_{0}$ and $\widetilde{X}_{0}$ decompose:
(1.5) $H^{k}(X_{0};\mathds{Q})\cong\underbrace{H^{k}(J\widetilde{C}_{0};\mathds{Q})\oplus H^{k-2}(J\widetilde{C}_{0};\mathds{Q}(-1))}_{\cong H^{k}(\widetilde{X}_{0};\mathds{Q})}\oplus H^{k-1}(J\widetilde{C}_{0};\mathds{Q}),$
@@
Notice that both are isomorphisms of mixed Hodge structures. The weight $k$ component of $H^{k}(X_{0};\mathds{Q})$ is given by $H^{k}(\widetilde{X}_{0};\mathds{Q})$, and the k1k-1 component is given by $H^{k-1}(J\widetilde{C}_{0};\mathds{Q})$. However, the decomposition (1.5) is non-canonical. On the other hand, since $\widetilde{X}_{0}$ is smooth and projective, $H^{k}(\widetilde{X}_{0};\mathds{Q})$ carries a pure Hodge structure, as is evident from the decomposition using the projective bundle formula in (1.6).
@@ 1.3.2. Abel Jacobi map on stable curves
In the last few decades, several authors have considered the problem of defining a similar map in the case of singular curves. The following is borrowed from [Ca et al.]. Let $C_{0}$ be a reduced, reducible, connected curve, and for each point $c\in C_{0}$, let $\mathcal{I}_{c}$ be its sheaf of ideals, which is torsion-free, rank 1, and of degree -1, and simple if $c\in C_{0}$ is a non-separating node. Then, if $C_{0}$ is free from separating nodes, and $x\in C_{0}$ is a smooth point,
@@ 1.3.3. Degenerations
    Following the constructions of Collino in [Co97, §1] among others, this assumption is somewhat natural.
@@ 2. Degeneration of curves and their Jacobians
In this section, we discuss degenerations of mixed Hodge structures of geometric origin. In particular, we focus on $H^{1}(C_{t};\mathds{Q})$ for families of smooth projective curves $C_{t}$, and we discuss the mixed Hodge structure we obtain in the limit of our families.
@@ 2.1. Survey of general degenerations of geometric origin
Let us set $\Delta\coloneqq\{z\in\mathds{C};\penalty 10000\ |z|<1\}$ and $\Delta^{\ast}\coloneqq\Delta\setminus\{0\}$. By Deligne, once $X_{0}$ is a complete simple normal crossing variety, then $H:=H^{*}(X_{0};\mathds{Q})$ can be equipped with a natural mixed Hodge structure coming from the Hodge structures provided by its normalization and by its irreducible components. On the other hand, Schmid constructs a mixed Hodge structure on $X_{0}$ using the information encoded in its smoothing. We now recall the main facts (see [Sch73]).
Let $f:\mathcal{X}\rightarrow\Delta$ be a semistable degeneration with central fibre $X_{0}$. Let the generic fibre $X_{t},t\in\Delta^{\ast}$ be smooth and of projective dimension $n$. Let $t_{0}$ be a smooth point, and $T:H^{*}(X_{t_{0}};\mathds{Q})\rightarrow H^{*}(X_{t_{0}};\mathds{Q})$ the Picard-Lefschetz transformation of the family. By Borel’s theorem, $T$ is a quasi-unipotent operator. Let $0<m\in\mathds{Z}$ be the smallest positive integer such that $T^{m}$ is unipotent. We define $N:=\frac{1}{m}\log T^{m}$ as the nilpotent logarithm of monodromy. By Landman’s theorem ([Lan73]), we have $N^{n+1}=0$. Let $\Gamma$ be the monodromy group of the family,

### 15 Apr 2026 v2

# Genus three Ceresa cycles and limit of archimedean heights

@@ Abstract.
For a one-parameter variation of biextension mixed Hodge structures, Brosnan and Pearlstein showed that the limit of the asymptotic height of the variation is given by a certain limit height of the nilpotent orbit. This limit height depends on the choice of a parameter. In the case of a variation of geometric origin related to Ceresa cycles associated with curves of genus three, after fixing a parameter, we show that this limit height is given by the Deligne splitting of a biextension mixed Hodge structure associated with cycles in the boundary.
@@
S. Goswami was partially supported by the Universitat de Barcelona María Zambrano postdoctoral fellowship. I. Spelta is a member of GNSAGA (INdAM). Both authors were partially supported by the Spanish MINECO research project PID2023-147642NB-I00.
@@ Introduction
Let $\Delta\coloneqq\{t\in\mathds{C}:\penalty 10000\ |t|<1\}$ be the open unit disc. Consider a family $\mathcal{C}\to\Delta$ of curves of fixed genus $g\geq 3$, with smooth and projective general fiber $C_{t},\penalty 10000\ t\neq 0$, and central fiber $C_{0}$ having nodal singularities. Two functions can naturally be attached to such a family. The first one is a holomorphic function $\nu(t)\coloneqq AJ(Z(C_{t}))$, known as a normal function and defined using the Abel-Jacobi image of the Ceresa cycle. Both $\nu$ and its derivative, the Griffiths infinitesimal invariant, have been studied widely by several authors (see, for example, [Gr83], [Voi88], [Ha90], [CP95],[Plz03], [GGK10], and [Ha25]). The second function is the height, which we introduce more generally. Let $X$ be a smooth complex projective variety, and let $Z$ and $W$ be homologically trivial algebraic cycles of codimensions $p$ and $q$, respectively, intersecting properly and satisfying $\dim X+1=p+q$. Under these assumptions, one necessarily has $|Z|\cap|W|=\emptyset$. To such a pair, one can associate a unique real number $\operatorname{ht}(Z,W)$, called the archimedean height. Hodge theoretically, $\operatorname{ht}(Z,W)$ measures how far a canonical subquotient $B_{Z,W}$ of the (rational) mixed Hodge structure $H^{2p-1}(X\setminus|Z|,|W|;\mathds{Q}(p))$ is from being $\mathds{R}$-split. $B_{Z,W}$ is an example of a biextension mixed Hodge structure with graded pieces
@@
A natural problem is to extend this construction from $\mathcal{M}_{g}$ to its Deligne–Mumford compactification $\overline{\mathcal{M}}_{g}$. It can be shown that while the biextension line bundle extends algebraically, the metric does not (see [Lear90], [HR04], [Pe06, Section 5]). In the language of heights, for a family of biextensions $\{B_{t}\}_{t\neq 0}$, instead of $\operatorname{ht}(B_{t})$, the asymptotic function $\operatorname{ht}(B_{t})+\mu\log|t|$ extends continuously to $t=0$, for $\mu\in\mathds{Q}$ that is controlled by the monodromy around $t=0$. This $\mu$ measures the singular behaviour of the height near the boundary. Then, the above study was generalized for an abstract biextension variation over a more general base in [BP19].
While the focus of the existing literature has been the study of $\mu$ and, more generally, the height jump divisor, the aim of our paper is slightly different. In [BP19, Definition 70], the authors introduce a limit height $H(N,F_{\infty},W)$ of the nilpotent orbit by ([BP19, Theorem 76])
$\lim_{t\to 0}\Big(\operatorname{ht}(B_{t})+\mu\log|t|\Big)=H(N,F_{\infty},W).$
Much like the limit mixed Hodge structure, the limit height depends on the choice of coordinate, while $\mu$ is independent of it. In this article, we aim to show that in a biextension variation of geometric origin, after fixing a choice of coordinate for $\Delta$, $H(N,F_{\infty},W)$ is ‘given by a biextension mixed Hodge structure attached to cycles in the boundary’. More precisely, following the approaches of [Co97, Ha90, Plz03], we consider the limit of asymptotic heights of Ceresa cycles associated to a special family of genus three curves as the first non-trivial case. We summarize our results as follows.
@@
Let $\mathcal{C}\to\Delta$ be a family of genus $g=3$ curves whose fiber $C_{t}$ is smooth and projective for $t\neq 0$, and whose central fiber $C_{0}$ is one of the following two types:
  • $C_{0}=D\cup E$, for two irreducible and smooth components $E$ and $D$ of genus one and two, respectively, with a separating node $y_{0}=D\cap E$.
Then the following happens:
@@
    When $C_{0}$ is irreducible and nodal, assuming 1.8, there is a choice of coordinate (assumption 2.2) such that the associated limit mixed Hodge structure $H^{1}_{\lim}(C_{t};\mathds{Q})$ is split. When $C_{0}$ is reducible, $H^{1}_{\lim}(C_{t};\mathds{Q})\cong H^{1}(D;\mathds{Q})\oplus H^{1}(E;\mathds{Q})$ is automatically split.
@@
    There exists a family of Ceresa cycles $\{Z(C_{t}),W(C_{t})\}_{t\in\Delta^{\ast}}$ that intersects properly in $JC_{t}$, such that the cycles $Z(C_{0})$ and $W(C_{0})$ also intersect properly in the compactification $X_{0}$ of the Jacobian $JC_{0}$, and are homologous to zero (Propositions 4.1 and 4.5).
  3. (iii)
    Under assumptions 1.8 and 2.2, the limit height $H(N,F_{\infty},W)$ associated with the biextension variation $B_{Z(C_{t}),W(C_{t})}$, is given by the Deligne splitting of the biextension mixed Hodge structure $B_{\widetilde{Z}(C_{0}),\widetilde{W}(C_{0})}$, where $\widetilde{Z}(C_{0})$ and $\widetilde{W}(C_{0})$ are the lifts of $Z(C_{0})$ and $W(C_{0})$ to the normalization $\widetilde{X}_{0}$ and are homologically trivial.
We mention that if $C_{0}$ is nodal and irreducible, there is a nontrivial vanishing cycle and a monodromy operator $N\neq 0$ with $N^{2}=0$ for $H^{1}(C_{t};\mathds{Q})$. The case where $C_{0}$ is reducible is simpler, since $N=0$ on $H^{1}(C_{t};\mathds{Q})$. Thus, in this case, the choice of coordinate is not explicit. Moreover, $JC_{0}$ is smooth and compact, so no normalization is needed in $(iii)$.
We now give an idea of the proof when $C_{0}$ is nodal and irreducible. Our method is inspired by the work in [Plz03] and [BP19], and can be thought of as a bridge between the two. Using assumption 1.8, we can fix a coordinate in such a way that the mixed Hodge structure $H^{1}_{\lim}(C_{t};\mathds{Q})$ is split (assumption 2.2). Then the key step involves finding the limit of the biextension variation
@@
The assumption 2.2 makes this study significantly easier. Nevertheless, the resulting limit mixed Hodge structure $\mathcal{V}_{\lim}$ no longer has the structure of a biextension. Consequently, the analysis of the limiting height amounts to extracting from $\mathcal{V}_{\lim}$ the graded pieces of the biextension $H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}(C_{0})|,|\widetilde{W}(C_{0})|;\mathds{Q}(2))$. This is the core motivation of our result (see Theorem 4.7), and is achieved using a purity result for the cohomology of ‘mildly’ singular varieties (see Proposition 4.3), and the injectivity of the Clemens-Schmidt morphism in our situation. Finally, we conclude by exploiting a relation between heights and certain morphisms of mixed Hodge structures. (see Proposition 3.4).
Remark.
Main Theorem(iii) is not merely an equality of two real numbers. The main point is that the limit height $H(N,F_{\infty},W)$ is obtained as a Deligne splitting of a geometrically defined biextension mixed Hodge structure appearing in the boundary. Heuristically, we are studying the interaction between the two mixed Hodge structures $H^{3}_{\lim}(JC_{t}\setminus|Z(C_{t})|,|W(C_{t})|;\mathds{Q}(2))$ and $H^{3}(\widetilde{X}_{0}\setminus|\widetilde{Z}(C_{0})|,|\widetilde{W}(C_{0})|;\mathds{Q}(2))$, which is much deeper than simply the equality of two invariants as real numbers.
The paper is organized as follows. In Section 1, we review basic material on mixed Hodge structures and on the Jacobian of a uni-nodal curve, and we introduce the two types of degenerations used throughout the paper. In Section 2, we recall the theory of limit mixed Hodge structures of geometric origin and apply it to analyze the mixed Hodge structure $H^{3}_{\lim}(JC_{t};\mathds{Q})$. In Section 3, we discuss biextensions, both abstractly and in the geometric setting, introduce the notion of height, and recall the variational aspects of [BP19]. Finally, in Section 4, we study families of Ceresa cycles, establish their properties and prove our main theorem (Theorem 4.7).
@@ 1. Preliminaries
In this section, we gather preliminary results from Hodge theory, nodal curves, and Jacobians.
@@ 1.1.1. Deligne splitting
Next, we show that $\delta$ commutes with morphisms of mixed Hodge structures.
@@ 1.2.1. Jacobian of a nodal curve
The semi-abelian variety $JC_{0}$ can be compactified. Indeed, identifying ${\rm Pic}^{0}(\widetilde{C}_{0})\cong{\rm Pic}^{0}(J\widetilde{C}_{0})$, the line bundle $L:=\mathcal{O}_{\widetilde{C}_{0}}(p-q)$ can be viewed as a line bundle over $J\widetilde{C}_{0}$, and $JC_{0}$ is the total space of $L$ minus the 0-section. Following [AK90], we give a quick description of the compactification: Let $Y_{0}$ and $Y_{\infty}$ be the zero and infinity sections of $\mathds{P}(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L)$ obtained through the projections $s_{0}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to\mathcal{O}_{J\widetilde{C}_{0}}$ and $s_{1}\colon\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\to L$ respectively. Then, the compactification $X_{0}=\overline{JC_{0}}$ is obtained as
@@
in other words the singular locus of $X_{0}$ is given by $J\widetilde{C}_{0}$, which is smooth and projective. Elements in $X_{0}$ are rank one torsion-free sheaves $M$ of degree 0 whose sections $s$ are allowed to approach the point $x_{0}$. The normalization $\widetilde{X}_{0}$ is given by the projective bundle $\mathds{P}\left(\mathcal{O}_{J\widetilde{C}_{0}}\oplus L\right)$ over $J\widetilde{C}_{0}$. From [UP24, Proposition 5.3] and the projective bundle formula for cohomology, the singular cohomology groups of $\widetilde{X}_{0}$ and $X_{0}$ decompose:
@@
(1.6) $H^{k}(X_{0};\mathds{Q})\cong H^{k}(\widetilde{X}_{0};\mathds{Q})\oplus H^{k-1}(J\widetilde{C}_{0};\mathds{Q}),$
Notice that both are isomorphisms of mixed Hodge structures. The weight $k$ component of $H^{k}(X_{0};\mathds{Q})$ is given by $H^{k}(\widetilde{X}_{0};\mathds{Q})$, and the k1k-1 component is given by $H^{k-1}(J\widetilde{C}_{0};\mathds{Q})$. However, the decomposition (1.6) is non-canonical. On the other hand, since $\widetilde{X}_{0}$ is smooth and projective, $H^{k}(\widetilde{X}_{0};\mathds{Q})$ carries a pure Hodge structure.
@@ 1.2.2. Abel Jacobi map on stable curves
In the last few decades, several authors have considered the problem of defining a similar map in the case of singular curves. The following is borrowed from [CCE08]. Let $C_{0}$ be a reduced, reducible, connected curve, and for each point $c\in C_{0}$, let $\mathcal{I}_{c}$ be its sheaf of ideals, which is torsion-free, rank 1, and of degree -1, and simple if $c\in C_{0}$ is a non-separating node. Then, if $C_{0}$ is free from separating nodes, and $x\in C_{0}$ is a smooth point,
@@ 1.2.3. Degenerations
    We remark here that even though this torsion condition is not very natural in general, the example we have in the back of our mind is the degeneration of a family of Ceresa cycles associated to a family of genus three curves, to the Collino cycle, as constructed in [Co97, Section (1.2)].
@@ 2. Degeneration of curves and their Jacobians
In this section, we discuss degenerations of mixed Hodge structures of geometric origin. In particular, we focus on $H^{1}(C_{t};\mathds{Q})$ for families of smooth projective curves $C_{t}$, and we discuss the mixed Hodge structure we obtain in the limit of our families. We fix the genus at $g=3$ to maintain parity with our main discussion, noting that similar results hold for general genera.
@@ 2.1. Survey of general degenerations of geometric origin
Let us set $\Delta\coloneqq\{z\in\mathds{C};\penalty 10000\ |z|<1\}$ and $\Delta^{\ast}\coloneqq\Delta\setminus\{0\}$. Further, let $f:\mathcal{X}\rightarrow\Delta$ be a semistable degeneration with central fibre $X_{0}$ which is a complete simple normal crossing divisor. Let the generic fibre $X_{t},t\in\Delta^{\ast}$ be smooth and of projective dimension $n$. By Deligne, once $X_{0}$ is a complete simple normal crossing variety, the $H:=H^{*}(X_{0};\mathds{Q})$ can be equipped with a natural mixed Hodge structure coming from the Hodge structures provided by its normalization and by its irreducible components. On the other hand, Schmid constructs a mixed Hodge structure on $X_{0}$ using the information encoded in its smoothing. We now recall the main facts (see [Sch73]).
Let $t_{0}$ be a smooth point, and $T:H^{*}(X_{t_{0}};\mathds{Q})\rightarrow H^{*}(X_{t_{0}};\mathds{Q})$ the Picard-Lefschetz transformation of the family. By Borel’s theorem, $T$ is a quasi-unipotent operator. Let $0<m\in\mathds{Z}$ be the smallest positive integer such that $T^{m}$ is unipotent. We define $N:=\frac{1}{m}\log T^{m}$ as the nilpotent logarithm of monodromy. By Landman’s theorem ([Lan73]), we have $N^{n+1}=0$. Let $\Gamma$ be the monodromy group of the family,